# Heterogeneous Treatment Effects and

### v3

@@ 4 HTEs and Mechanisms
However, postulating a MDV is not sufficient to use HTEs to detect the activation of a mechanism. Specifically, we must also be concerned with whether a given covariate is a MDV for other mechanisms (and therefore incorporated in the direct effect). To this end, researchers need an exclusion restriction to limit the number of mechanisms that are moderated by a given covariate, $X_{k}$. Logically, if a single covariate moderates a mechanism of interest and the direct effect, a finding of HTEs in that covariate cannot isolate the mechanism of interest. Assumption 1 rules out the possibility that direct effects depend on $X_{k}$. In the more general multiple-mechanism case, we also need exclusion assumptions to rule out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$ (see Appendix C).
Assumption 1 (Exclusion).
Given $z,z^{\prime}\in Z$ and $x,x^{\prime}\in X_{k}$, $X_{k}$ is excluded to the direct effect such that $ADE(z,z^{\prime};X_{k}=x)=ADE(z,z^{\prime};X_{k}=x^{\prime})$.
Assumption 1 constrains the relationship between a moderator, $X_{k}$, and the direct effect (taken to be a composite of other possible mechanims). Figure 3 depicts a violation of Assumption 1 graphically for a MDV ($X_{k}$) of the mechanism $M$. Assumption 1 rules out the blue dashed path from $X_{k}$ to the direct effect of $Z$ on $Y$. Given our notation for interactions in a DAG, this means that $X_{k}$ cannot moderate the direct effect of $Z$ on $Y$. In the absence of Assumption 1, we cannot link a heterogeneous treatment effects to a unique theoretical mechanism representation (here, $M$) because we have no way to ensure additive separability of the effect of $M$ from the effects of other mechanisms. Importantly, Assumption 1 does not imply that $X_{k}$ must be independent of the direct effect of $Z$ on $Y$; these quantities can be arbitraily correlated so long as $X_{k}$ does not induce heterogeneity in the direct effect.
$Z$$M$$Y$$X_{k}$
Figure 3: Assumption 1 rules out the blue dashed path. All black and orange paths are permissible under Assumption 1. For $X_{k}$ to be a MDV, at least one of the orange paths must be present.
Use of HTEs in order to learn about mechanism activation represents one alternative to mediation analysis. Assumptions 1 forms the core assumption underpinning this use of HTE. The comparison to mediation analysis invites a comparison of the exclusion assumption to the assumption of sequential ignorability. It is useful to note that there is no logical ordering of the two types of assumptions: the exclusion assumption does not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumption. This means that HTEs cannot said to be a “more agnostic” or “less agnostic” quantitative test of mechanism activation than mediation. [^14] Mediation analysis attempts to estimate the influence of different mechanisms by decomposing total effects into other causal estimands. When our primary interest is the activation of a mechanism (as in HTEs analysis), we can say that a mechanism is activated in mediation analysis if its indirect effect is non-zero. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in Appendix E.
4.1 HTEs and Mechanisms with Directly-Affected Outcomes
Collectively, the concept of MDVs and the exclusion assumption convey the basic intuition that our initial results draw upon. For HTEs to provide information about a mechanism, it must be the case that the moderator of interest affects: (1) the degree to which treatment activates a mechanism or (2) the mechanism’s effect on our outcome of interest, $Y$. Recognition of the former scenario—consistent with our stylized example of voter updating—is well-known. The latter opens new possibilities for identifying moderators. The exclusion assumption suggests that if a covariate, $X_{k}$, is a MDV for direct effects (or multiple mechanisms), it cannot be used to confirm the presence of a given mechanism. This logic is straightforward, but it remains implicit in most uses of heterogeneous treatment effects to detect mechanisms. These intuitions give rise to Proposition 1.
Proposition 1.
Suppose that $Y$ is directly affected by mechanism $M$ and Assumption 1 holds with respect to $X_{k}$. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M$.
Proposition 1 conveys important implications about our ability to use heterogeneous treatment effects to provide evidence for a mechanism. Recall that if $X_{k}$ is a MDV for mechanism $M$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for some $x,x^{\prime}\in X_{k}$. This provides evidence that mechanism $M$ is active for at least one unit. The exclusion assumption rules out the possibility that $ADE(X_{k}=x)\neq ADE(X_{k}=x^{\prime})$. When the assumption holds, HTEs in $X_{k}$ are thus sufficient to show that $X_{k}$ is a MDV for mechanism $M$. This conforms to standard interpretations that the presence of HTEs support arguments about mechanism activation. Nevertheless, this interpretation invokes the exclusion assumption, which is generally not invoked explicitly. In our running example, we showed the existence HTE in the voters’ priors ($\pi_{i}$) for the directly-affected voter utility outcome. If one were to detect heterogeneity in an empirical study in this class, the heterogeneity would provide evidence that voter learning is an active mechanism.
We now consider the converse: the case when there exist no HTEs with respect to $X_{k}$.
Proposition 2.
Suppose that $Y$ is directly affected by mechanism $M$ and Assumptions 1 and 2 hold. If no HTEs exist with respect to $X_{k}$, at least one of the following must be true:
  1. 1.
    $X_{k}\notin\textbf{X}^{MDV}$ for mechanism $M$.
  2. 2.
    No MDV exists.
Proposition 2 shows that a lack of HTEs provides less information with regard to mechanism activation than is generally asserted. Under the exclusion assumption, there are two reasons why HTEs may not exist with respect to a covariate, $X_{k}$. First, it may be the case that $X_{k}$ is not a MDV for mechanism $M$. In this sense, we have misspecified the theoretical relationship between a given covariate and a mechanism. Second, it may be the case that no MDV exists for mechanism $M$. As we discuss in Corollary 1, there are two possible reasons why a MDV would not exist for mechanism $M$. Importantly, we show that this could happen with an active or an inert mechanism $M$.
Corollary 1.
If no MDV exists for a mechanism $M$, there are two possibilities:
(1) Mechanism $M$ is not active.
(2) Mechanism $M$ is active, but there exists no $X$ for which $AIE(X=x)\neq AIE(X=x^{\prime})$.
Case (1) of Corollary 1 is implied by the definition of MDV. If a mechanism is inert—thereby producing an indirect effect of zero for all units—there cannot exist any MDVs, measured or unmeasured. In contrast, in Case (2), a mechanism can be active and produce the same indirect effect for all units. In this case, there are no covariates that moderate the indirect effect. These results show that, in contrast to standard interpretation, a lack of heterogeneity cannot tell us about whether a mechanism is active. Moreover, our theory could be misspecified, meaning that our postulated MDV, $X_{k}$ is not actually a MDV. An assessment of HTEs with respect to a single moderator cannot distinguish between these three possibilities. Nor can we assign probabilities to these (non-mutually exclusive) explanations.
In our running example, recall that theoretically there do not exist HTEs in valence for the directly-affected voter utility outcome. Within the model, we know that the valence mechanism is inert. However, in an empirical exercise, the analogous finding of no detectable heterogeneity provides limited information. Proposition 2 and Corollary 1 enumerate three possibilities: we chose the wrong MDV for an active valence mechanism; the valence mechanism is inert (as in the theory); or there exists no MDV for an active valence mechanism. Crucially, the empirical finding of no (detectable) heterogeneity cannot distinguish between these possibilities. Therefore, we cannot “rule out” a mechanism due to a lack of observed HTEs.
Comparing Propositions 1 and 2, the presence of HTEs provides more information with regard to mechanisms than does the absence of HTEs. In this sense, relying upon the a lack of heterogeneity to “rule out” a potential mechanism would require even stronger assumptions. Specifically, we would need to assume that $X_{k}\in\textbf{X}^{MDV}$ in order to rule out the activation of mechanism $M$. Indeed, such an assumption is precisely what we are trying to learn from the presence of HTEs in Proposition 1.
5 Indirectly-Affected Outcomes
When outcomes are directly affected by a mechanism under a theoretical model, the existence of HTEs provides evidence that a mechanism is active, as we have shown in Proposition 1. Yet, the situation is more complicated in the case of indirectly-affected outcomes. As in our discussion of the exogenous shocks and voting, for the vote choice outcome, there exist HTEs in both prior beliefs (the mechanism in the model) and in valence (not the mechanism in the model). In this section, we show that this finding is general to situations with indirectly-affected outcomes as defined in Definition 3.
This represents a large class of mappings that are used in most empirical applications. It includes discrete choices made on the basis of comparisons in utility as in the theoretical motivating example. It also applies to mappings from an attitude to a $Q$-item Likert scale of the form:
$h(Y)$ =$=\begin{cases}1&Y\in(-\infty,c_{1}]\\ 2&Y\in(c_{1},c_{2}]\\ \vdots&\\ Q&Y\in(c_{Q-1},\infty),\end{cases}$ (15)
in which $c_{t}$ are increasing thresholds in a latent attitude. Our focus for this section is outcomes that are indirectly affected such that they are produced by a non-linear mapping from a directly affected outcome.
To understand what HTEs reveal with respect to an indirectly-affected outcome, it is useful to introduce one final concept. We will denote $\textbf{X}^{R}$ as the set of all possible pre-treatment covariates with non-zero effects on the directly-affected outcome, $Y$ [^15] Formally, if $X_{k}\in\textbf{X}^{R}$, then there exist $x\neq x^{\prime}\in X_{k}$ such that $Y(Z,M(Z;X_{k}=x,X_{\neg k});X_{k}=x,X_{\neg k})\neq Y(Z,M(Z;X_{k}=x^{\prime},X_{\neg k});X_{k}=x^{\prime},X_{\neg k})$. Covariates in $\textbf{X}^{R}$ can be thought of as “relevant” for predicting outcome $Y$. It is also useful to let X be the set of all possible pre-treatment covariates. It is clear that for any outcome, $Y$, and mechanism $M$, $^{MDV}\subseteq\textbf{X}^{R}\subseteq\textbf{X}$ X. Typically, these subsets will be proper.
We now return to our main question of interest: what do HTEs reveal with regard to mechanisms? Proposition 3 considers the case when there are HTEs in a covariate $X_{k}$. Here, we can learn that $X_{k}\in\textbf{X}^{R}$, but this is not informative about whether $X\in\textbf{X}^{MDV}$, since $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$. In order to make an inference about mechanism $M$ (under the exclusion assumption) we need to know whether $X_{k}\in\textbf{X}^{MDV}$. If $X_{k}\in\textbf{X}^{MDV}$ for mechanism $M$, then mechanism $M$ is active. If $X_{k}\notin\textbf{X}^{MDV}$, mechanism $M$ may or may not be active. The intuition for this result is straightforward. The non-linear mapping from $Y$ to $h(Y)$ “breaks” the additive separability between any $X_{k}\in\textbf{X}^{R}-\textbf{X}^{MDV}$ and the indirect effect.
Proposition 3.
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumptions 1holds. If HTEs exist with respect to $X_{k}$, then $X_{k}\in\textbf{X}^{R}.$
The result and logic in Proposition 3 shows precisely why we see HTEs in both the priors and valence with the vote choice outcome in our running example. Even though valence is unrelated to the learning mechanism, it does enter the voter’s utility in (2). This means that a variable measuring valence would be relevant for the the directly-affected outcome. Yet, because vote choice is a non-linear function of this outcome, there exist—and we may observe—HTEs with respect to valence. This suggests that the qualitative existence HTEs with respect to an indirectly-affected outcome are not useful for the detection of mechanisms.
It is useful to consider a simple numerical example of Proposition 3 where $X_{k}\in\textbf{X}^{R}-\textbf{X}^{MDV}$, meaning that $X_{k}\in\textbf{X}^{R}$ but $X_{k}\notin\textbf{X}^{MDV}$. Suppose that we are interested in how a mobilization treatment, $Z_{i}\in\{0,1\}$, affects citizens’ decision to vote. Consider two covariates that predict turnout: $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\in\{0,1\}$. We will further assume that $X_{1}$ is a MDV for the unique mechanism, such that:
$M(Z;\textbf{X})$ $=(1+Z)X_{1}$
Potential voters’ utility from voting is given by:
$U(Z,X)$ $=M(Z;X_{1})+X_{2}=(1+Z)X_{1}+X_{2}$
$h(U(Z,X))$ =$=\begin{cases}1&\text{ if }(1+Z)X_{1}+X_{2}\geq 0\\ 0&\text{ else }\end{cases}$
In this case, Assumption 1 holds by construction since there is only one mechanism. Now, suppose that a researcher mistakenly thought that $X_{2}$ was a MDV for a mechanism (either $M$ or an inert mechanism). Since we have constructed the data generating process, we know that it is not: $X_{2}\in\textbf{X}^{R}-\textbf{X}^{MDV}$. Evaluating the CATE of $Z$ on turnout when $X_{2}=1$, we have:
$CATE(X_{2}=1)$ $=E[h(U(Z=1,X))-h(U(Z=0,X))|X_{2}=1]$
$=\Pr(2X_{1}+1>0)-\Pr(X_{1}+1>0)$
$=\Phi(-1)-\Phi(-\frac{1}{2})$
$\approx-0.15$
While it is straightforward to construct infinte such examples numerically, it is reasonable to ask: is this plausible in “real world” examples? Unfortunately, it is not possible to provide a purely empirical example because mechanisms are theoretical, not empirical objects. While we can, in principle, observe and measure the effects of a mechanism, we do not observe mechanisms themselves (Slough and Tyson, 2024a). [^16] The status of mechanisms as theoretical, not empirical, objects is precisely why mechanism testing is so difficult. However, simulations using real data and an assumed theory allow us to assess the likelihood that we observe heterogeneity when a covariate is or is not an MDV. In Appendix H, we use 2020 ANES data in a Monte Carlo simulation motivated by the theoretical model of Little, Schnakenberg, and Turner (2022). This paper proposes two mechanisms that account for how citizens update beliefs in response to new information: accuracy and directional motives. Their model implies that partisanship (or ideology) should be a MDV for directional but not accuracy motives, which would allow researchers to look for heterogeneity in partisanship to assess the presence of directional motives as a mechanism.
Following this logic, we simulate different treatment effects on a measure of latent attitudes about greenhouse gas regulation. Some simulations allow for directional motives, others shut down this channel. We show that when treatment has a non-zero effect on the latent attitudes for any subset of partisans, as we would expect from accuracy motives alone, there are HTEs in partisanship on a binary measure of preferences for stronger greenhouse gas regulation. This means that we cannot use these HTEs to assess the presence of directional motives with this indirectly measured outcome. Further, using the ANES sample, for some simulated effect sizes, there is greater statistical power to detect heterogeneity in partisanship when simulated effects on the latent variable are homogeneous in partisanship than when they are heterogeneous.
We now return to a final case of our theoretical analysis by asking when we are examining an outcome that may be indirectly affected by mechanism $M$, what can we learn from a lack of HTEs? Proposition 4 indicates that in this case, we can infer that $_{k}\in\textbf{X}$ X. This is obviously a vacuous result. We already know that $_{k}\in\textbf{X}$ X since $X_{k}$ is a covariate and X is the set of all covariates. We purposely state a vacuous result to emphasize how little can be ascertained about mechanisms from the non-existence of HTEs when outcomes are indirectly affected by a mechanism.
Proposition 4.
Suppose that observed outcome $h(Y)$ is a non-linear mapping of directly-affected outcome $Y$ and Assumption 1 holds. If HTEs do not exist with respect to $X_{k}$, then $_{k}\in\textbf{X}.$ X .
Proof.
The result follows simply because if $_{k}\notin\textbf{X}$ X, then $X_{k}=\emptyset$. ∎
Often we make assumptions about the mapping $h$. For example, the mapping in (15) imposes assumptions about how latent attitudes translate into Likert-scale responses. When we are willing to make such assumptions, we can refine Proposition 4 slightly. Specifically, in Proposition 9, we show that under the exclusion assumption, for absolutely continuous directly-affected outcome, $Y$ and the non-linear transformation in (15) (for any $Q\geq 2$ categories), if HTEs do not exist with respect to $X_{k}$, then $X_{k}\notin\textbf{X}^{R}$. Because $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$ we know then that $X_{k}\notin\textbf{X}^{MDV}$ if $X_{k}\notin\textbf{X}^{R}$. But as in Proposition 2 and Corollary 1, there are multiple possible explanations: our theory about how $X_{k}$ relates to mechanism $M$ could be wrong or no MDV exists for mechanism $M$. These possibilities mean that we cannot make an inference about mechanism (non)-activation from the absence of HTEs with respect to $X_{k}$.
In sum, our propositions characterize four cases into which we can classify attempts to ascertain mechanism activation from HTEs, as described in Table 2. On the columns, we stratify by whether the outcome is directly affected by the mechanism or whether it is indirectly affected (via some non-linear transformation of the directly-affected outcome). On the rows, we consider whether there exist HTEs in a covariate of interest, $X_{k}$. Our results show that, under exclusion assumptions, this strategy provides information about mechanism activation in one case: when HTEs exist for a directly-affected outcome. In the other cases, HTEs provide incomplete—or no—information about the activation of a mechanism of interest.
Outcome variable is:
Directly affected by mechanism $M$ Indirectly affected by mechanism $M$
$\exists$ HTEs wrt $X_{k}$: $X_{k}\in\textbf{X}^{MDV}$ $X_{k}\in\textbf{X}^{R}$
$\implies M$ is active. $M$ is active or inactive
$\nexists$ HTEs wrt $X_{k}$: $X_{k}\notin\textbf{X}^{MDV}$ and/or $\textbf{X}^{MDV}=\emptyset$ $_{k}\in\textbf{X}$ X (vacuous)
$M$ is active or inactive $M$ is active or inactive
Table 2: In this table, $\textbf{X}^{MDV}$ is defined with respect to mechanism $M$. The results in each cell invoke Assumption 1.
6 Implications for Applied Research
Our framework and analysis holds a number of implications for applied research that seeks to study causal mechanisms using HTE. We discuss implications and recommendations in four categories:
  1. 1.
    Desiderata for applied theory in empirical research
  2. 2.
    Improvements in the interpretation of HTE
  3. 3.
    Recommendations for prospective research design
  4. 4.
    Benefits and limitations of stronger theoretical assumptions
6.1 Three essential theoretical questions
Our framework suggests that some form of theory is necessary to link HTEs to a causal mechanism. While our analysis is ultimately agnostic with respect to the type of theory (e.g., formal or informal), our framework lays out three attributes of a theory that are needed to support any analysis of causal mechanisms using HTE.
  1. 1.
    A set of candidate mechanisms. Researchers must generate a list of the candidate mechanisms that may mediate the effect of $Z$ on $Y$.
  2. 2.
    The relationship between a covariate, $X_{k}$, and each candidate mechanism. This requires answering two questions:
    1. (a)
      For which mechanism, $M$, is $X_{k}$ a candidate MDV?
    2. (b)
      Is Assumption 1 plausible?
  3. 3.
    Specifying the relationship between the theoretical outcome of interest and measured outcomes. Which outcome(s) are directly affected by the candidate mechanisms? Which are indirectly affected by those outcomes?
Question #1 is fairly standard in applied empirical research. Researchers often posit one or more mechanisms of interest in addition to alternative explanations. In the exposition of our framework, we focus on one mechanism and relegate alternative explanations to the composite direct effect. Appendix C provides guidance for cases with multiple mechanisms of interest.
Questions #2 and #3, instead, are much less standard. When researchers assess HTEs, it is rare to discuss the relationship between the moderator of interest and other mechanisms (within the direct effect) even though this assumption is necessary (but not sufficient) for mechanism attribution with HTEs, as we show. Explicit justification of a moderator of interest as a candidate MDV of a single mechanism may facilitate the search for MDVs. As we show in Figure 2, a MDV can moderate the effect of treatment on the mediator or the effect of the mediator on the outcome. Since the latter possibility is generally ignored in applied research, these considerations may broaden the set of candidate MDVs.
Question #3 is not a standard consideration in applied research. We typically do not distinguish between directly-affected outcomes ($Y$’s) and indirectly-affected outcomes ($h(Y)$’s). Yet, as we have shown, this distinction is critical to the ability of HTEs to provide information about causal mechanisms. In general, formal theoretic treatments permit straightforward evaluation of whether $h(\cdot)$ is a non-zero linear function, which can tell us whether HTEs could be informative about mechanisms. More broadly, however, Question #3 shows that we may want to evaluate HTEs for some outcomes but not others, and also provides a principled justification for this determination.

### v4

@@ 4 Applications and Guidance for Research Design
(a) Distaste for corruption ($M$) and partisan alignment ($a$) are distinct and additively separable arguments in voters’ preferences for the incumbent ($y_{1}$), as in the running example.
Source: Eggers (2014)
[…]
(b) There are two mechanisms, distaste for corruption ($M$) and voter coordination motives ($R$). Partisan alignment ($a$) is distinct from both mechanisms and additively separable arguments in voters’ preferences for the incumbent ($y_{1}$), as in the running example.
Source: Arias et al. (2019)
[…]
(c) Partisan alignment (bias) affects voter processing of corruption information via motivated reasoning and enters voters’ preferences for the incumbent through a distinct channel.
Source: Anduiza et al. (2013)
[…]
(d) Partisan alignment ($a$) conditions corruption aversion ($\lambda$), which moderates distaste for corruption ($M$). Alignment also enters voter’s preferences for the incumbent ($y_{1}$) through a distinct channel.
Source: de Figueiredo et al. (2023)
Figure 3: Four theoretical accounts of how partisan alignment (or bias) and information relate to voter preferences for the incumbent and vote choice. $c$ refers to corruption information; $\lambda$ is a voter’s corruption aversion; $M$ is voters’ distaste for observed corruption; $a$ is voters’ partisan bias toward the incumbent; $y_{1}$ is voter utility (preference) for the incumbent; and $y_{2}$ is vote choice for the incumbent.
Our framework suggests four avenues for improved use of HTEs for mechanism detection which we explore below. First, it suggests a set of necessary theoretical considerations/arguments. Second, it suggests improvements in the interpretation of observed HTEs. Third, it offers recommendations for prospective research design. Finally, it provides guidance on when stronger theoretical assumptions may be useful to impose.
4.1 Three essential theoretical considerations
Our framework identifies three attributes of a theory that are needed to support any analysis of causal mechanisms using HTE. First, researchers must identify a set of candidate mechanisms. The examples we cite are clear in positing a set of candidate mechanisms. Three—Panels (a), (c), and (d)—focus on a single mechanism: distaste for an incumbent’s corruption. One—Panel (b)—instead suggests that distaste and voter coordination motives are two distinct mechanisms through which corruption information affects voter utility. The applications we discuss are quite clear about the mechanisms thought to underlie the effects that they measure.
Second, in order to use HTE to learn about a given mechanism, however, our analysis shows that one must (1) invoke exclusion assumptions and (2) identify candidate MDVs. In theories with a single mechanism—as in Panels (a), (c), and (d) of Figure 3—it is only necessary to defend Assumption 1 for a given covariate. In the case of the partisan alignment covariate, this means that we should justify the absence of a direct effect of corruption information that varies in partisan alignment. In Panel (b), when there are multiple mechanisms, one would need to invoke Assumptions 1 and 2 for a given covariate, e.g., partisan alignment. As represented in the papers (and therefore in Figure 3), the exclusion assumptions should hold for the voter utility outcome ($y_{1}$) in each case. We note that because these studies focus on a small (and well-articulated) set of mechanisms, assessing whether Assumptions 1 and (where relevant) 2 are plausible is relatively straightforward. In other types of studies that propose a larger set of candidate mechanisms, (1) additional exclusion assumptions must be invoked for each candidate mechanism; and (2) justification of these exclusion assumptions becomes more difficult because there are more plausible violations of these assumptions.
The different theories vary in whether partisan alignment, $a$, should serve as an MDV for voter utility, $y_{1}$. In the first two theories, depicted in Panels (a) and (b), partisan alignment is not a candidate MDV. We can see this because there is no arrow from $a$ to the edge between $c$ and $M$ or the edge between $M$ and $y_{1}$, or (in Panel (b)), the voter coordination mechanism ($R$). In contrast, $a$ serves as an MDV for the distaste mechanism in Panels (c) and (d). We can see this from the direct arrow from $a$ to the edge between $c$ and $M$ in Panel (c)—which captures voters’ motivated reasoning about the corruption disclosure—and from $a$ to $\lambda$ to the edge between $c$ and $M$ in Panel (d), which captures the idea that partisan affiliation shapes corruption aversion, which in turn conditions the degree to which corruption information ($c$) gnerates a distaste for observed corruption.
Third, authors must consider the relationship between outputs of the candidate mechanisms and measured outcomes. The empirical studies associated with Panels (a), (b), and the field experiment in (d) measure aggregate vote share at the level of the constituency (Panel a) or precinct (Panels b and d). Within our framework, aggregate vote share can be thought of as $\sum_{i}y_{i2}/n$, where $n$ is the number of voters. [^18] In Panel (a), Eggers (2014) also provides a survey-based measure of $y_{2}$ from the British Election Survey by analyzing self-reported vote choice. Some studies also consider vote share as a fraction of registered voters. This would require a model that allows voters to abstain, but yields substantively similar results. This means that we should expect HTEs in partisan alignment for each of these studies. But recall that of these three studies, partisan alignment is only an MDV under the theory in (d) in which alignment conditions a voter’s corruption aversion. In contrast, the survey-based studies in Panel (c) and the survey experiment component of Panel (d) measure outcomes that resemble voter utility through Likert scales. In both relevant Panels, the authors expect partisan affiliation to moderate the effect of corruption information on distaste for corruption (albeit for different reasons). So in contrast to the our model (and the examples in Panels (a) and (b)), we would expect to observe HTE in partisan affiliation for these outcomes if the distaste for corruption mechanism is active.